OpenStudy (anonymous):

Integrate dx/(25-16x^2)

OpenStudy (anonymous):

No one?

OpenStudy (anonymous):

Never mind, I made trough the question. :)

OpenStudy (apoorvk):

\[\large\int\frac{dx}{a^2 - (x)^2} = \frac 1 {2a}ln(\frac{x+a}{x-a}) \] Can/did you use this? Sorry that I took so long to type a reply.

OpenStudy (apoorvk):

'+ c' --> constant of integration, ofcourse.

OpenStudy (anonymous):

I'll turn on my pc to answer you, wait a sec

OpenStudy (anonymous):

try this..\[4x=5sinu\]

OpenStudy (anonymous):

Nope, can't use that. I used Integration by substitution: First i had to ajust my equation: \[\int\limits_{?}^{?} dx/\sqrt{25-16x ^{2}} = 1/5\int\limits_{?}^{?} dx/\sqrt{1-16x^{2}/25}\] Ready to go: \[1/5\int\limits_{}^{}dx/\sqrt{1-16x ^{2}/25}\] \[u=4x/5\] \[du=4/5\] \[1/4\int\limits\limits_{}^{} dx/\sqrt{1-u ^{2}}\] \[\int\limits\limits_{}^{} dx/\sqrt{25-16x ^{2}} = (1/4)\sin ^{-}(4x/5) + C\]

OpenStudy (anonymous):

I solved it in that way..

OpenStudy (anonymous):